OpenStudy (anonymous):

1/r^8+r^0.4 find ds/dt

OpenStudy (abb0t):

\(\large \frac{ds}{dt}\)? If you're looking for\(\large \frac{ds}{dt}\) then your answer is 0

OpenStudy (anonymous):

s=1/r^8+r^0.4 find ds/dt

OpenStudy (valpey):

Do you know anything about the relationship between r and time? Perhpas r is a rate like a speed where position (s) changes over time (t)?

OpenStudy (anonymous):

I just know I need to find the derivative

OpenStudy (valpey):

If r does not change with respect to t then the answer is zero; but I would make sure you know that.

OpenStudy (anonymous):

ok thank you

OpenStudy (abb0t):

Do you even lift?